OpenStudy (anonymous):

A seller buys an item from a manufacturer for $9,500. He increases the cost by 5 percent. What is the markup value? markup = original cost × percentage markup

OpenStudy (anonymous):

9500*1,05. Or did i misunderstand the question?

OpenStudy (jiteshmeghwal9):

\[\huge{9,500+{5\over100} \times 9500}\]

OpenStudy (jiteshmeghwal9):

right ???

OpenStudy (jiteshmeghwal9):

solve this @lilhen & u ill gt ur answer :)

OpenStudy (anonymous):

ok

OpenStudy (mayankdevnani):

no @jiteshmeghwal9 you are wrong

OpenStudy (anonymous):

nope

OpenStudy (mayankdevnani):

\[9500\times \frac{5}{100}\]

OpenStudy (jiteshmeghwal9):

why ?

OpenStudy (mayankdevnani):

and you get RS475

OpenStudy (mayankdevnani):

because markup = original cost × percentage markup

OpenStudy (mayankdevnani):

ok @jiteshmeghwal9

OpenStudy (mayankdevnani):

and markup value=Rs475 ok @lilhen

OpenStudy (anonymous):

looks good

OpenStudy (jiteshmeghwal9):

ohh ! srry i gave u the selling price really sorry :(

OpenStudy (mayankdevnani):

yw it's right

OpenStudy (anonymous):

its alright

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

it happens to us all

OpenStudy (mayankdevnani):

but GOOD JOB @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

thanx !

OpenStudy (mayankdevnani):

welcome